Can you solve this please?
2x+2y=2
3x+y=1
Can you use the substituion Method?
I have NO idea.

Answer by stanbon(48558)

Solve the 2nd equation as follows: y=-3x+1
Substitute that into the 1st equation to get:
2x+2(-3x+1)=2
Divide thru by "2" to get:
x+(-3x+1)=1
-2x+1 = 1
-2x=0
x=0
Put that value of "x" into either of the original equations, as follows:
0+2y=2
y=1
Check those values for "x" and "y" in the other original equations:
3(0)+1=1
1=1
So the solution for the system is x=0, y=1.
Cheers,
Stan H.